OpenStudy (anonymous):

What is the slope of the line whose equation is -48 = 2x -8y?

OpenStudy (anonymous):



OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

lastttt question i think the answer is 6/24

OpenStudy (sleepyjess):

First add 8y to each side

OpenStudy (sleepyjess):

8y-48=2x now add 48 to both sides

OpenStudy (anonymous):

sorry okay so i'm writing down on the paper 8y-48=2x

OpenStudy (anonymous):

and i add -48 to both sides right?

OpenStudy (anonymous):

and then what do i do

OpenStudy (sleepyjess):

no just add +48

OpenStudy (anonymous):

oh ya you change the sighns right?

OpenStudy (anonymous):

do i bring everything else down

OpenStudy (anonymous):

or do i add 48 + 48

OpenStudy (sleepyjess):

8y-48+48=2x+48

OpenStudy (anonymous):

@sleepyjess you there

OpenStudy (anonymous):

oh okay then i'd get hold on

OpenStudy (anonymous):

8y - 96 = 50x

OpenStudy (anonymous):

am i wrong?

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

just 8y=2x+48

OpenStudy (anonymous):

but there looking for a fraction

OpenStudy (anonymous):

nvm i'll just close this and get out of your hair

OpenStudy (sleepyjess):

yes, now divide by 8 on each side to get y=2/8x+6

OpenStudy (sleepyjess):

and 2/8 simplifies to 1/4